# Bayesian Shrinkage in High-Dimensional VAR Models: A Comparative Study Harrison Katz1 and Robert E. Weiss 2

### v2

@@ 1 Introduction
Vector autoregressive (VAR) models remain a cornerstone of multivariate time‑series analysis Sims, (1980); Chan, (2020); Koop and Korobilis, (2013). A dd‑dimensional VAR($p$) posits that an observed series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ satisfies
@@
where $\boldsymbol{\varepsilon}_{t}\!\sim\!\mathcal{N}(\mathbf{0},\Sigma_{\varepsilon})$ is white noise. VARs are ubiquitous in macroeconomics and finance; see Stock and Watson, (2002); Crump et al., (2021); Carriero et al., (2022); Zhou and Chan, (2023) for recent surveys and methodological refinements.
@@
Because a VAR($p$) contains $d^{2}p$ coefficients, dimensionality rises quickly with either $d$ or $p$ Bańbura et al., (2010); Koop, (2013); Korobilis and Pettenuzzo, (2019). Shrinkage—penalising or down‑weighting coefficients toward simpler structures—therefore plays a critical role in stabilising estimation and forecasts Stock and Watson, (2002); Chan, (2021), especially when $d^{2}p\gg T$ Bai et al., (2022).
@@
Bayesian shrinkage. Local–global priors such as the horseshoe Carvalho et al., (2010) or Bayesian lasso Park and Casella, (2008) shrink most coefficients strongly while allowing a few to escape toward their likelihood values. Variants include the spike‑and‑slab George and McCulloch, (1997), hierarchical horseshoes Makalic and Schmidt, (2016); Prüser, (2021), order‑invariant priors Chan et al., (2024), dynamic sparsification for hybrid TVP‑VARs Chan, (2023), and structured factor‑augmented horseshoes Zhou and Chan, (2023). Semi‑global or block‑specific shrinkage extends the idea by letting related groups share a common scale parameter Gruber and Kastner, (2025); Prüser and Blagov, (2022). Computational advances now make these priors feasible in very large systems: fast variational Bayes Bernardi et al., (2024), approximate Bayes for huge multi‑country VARs Huber et al., (2022), and conjugate subspace shrinkage Huber and Koop, (2023). Pandemic‑robust “outlier‑aware” priors improve real‑time performance during extreme events Cascaldi-Garcia, (2022). Other recent Bayesian shrinkage methods in time series have been explored Huber and Koop, (2023); Kowal et al., (2019); Katz et al., (2024, 2025).
@@
Frequentist and machine‑learning regularisation. Ridge regression Doan et al., (1984) remains popular, while equation‑by‑equation Lasso or elastic‑net penalties Tibshirani, (1996); Zou and Hastie, (2005); Sánchez García and Cruz Rambaud, (2022) yield sparse high‑dimensional VARs with frequentist guarantees under weak dependence Masini et al., (2022). Envelope and reduced‑rank ideas provide additional dimension‑reduction tools Samadi and Herath, (2023); Cubadda and Hecq, (2022), and non‑linear factor compression has recently been explored Klieber, (2024). Structured penalties—group Lasso, hierarchical lag shrinkage, SCAD, MCP—continue to improve point forecasts and interpretability Nicholson et al., (2017); Basu et al., (2019); Nicholson, (2020); Song and Bickel, (2019); Chen and Chen, (2021). New inferential results ensure valid impulse‑response analysis even after sparse estimation Krampe et al., (2023).
@@
Forecasting and empirical performance. Time‑varying parameter BVARs with horseshoe or normal‑gamma shrinkage adapt well to structural change Bitto and Frühwirth-Schnatter, (2019); Feldkircher et al., (2024). Factor‑adjusted network VARs (FNETS) separate common and idiosyncratic dynamics to enhance both forecasts and network interpretation Barigozzi et al., (2024). Large‑scale empirical comparisons consistently find that local–global or semi‑global priors outperform purely global shrinkage and many frequentist alternatives in density and point forecasts Huber and Feldkircher, (2019); Aprigliano, (2020); Gefang et al., (2023).
@@ 2.1 Bayesian Shrinkage Priors
where $\mathbf{L}_{\Omega}$ is the Cholesky factor of a correlation matrix with an LKJ prior Lewandowski et al., (2009), and each component of $\sigma$ follows a half-Cauchy prior. This flexible structure permits correlation among the $d$ error components.
@@
The horseshoe prior Carvalho et al., (2010); Makalic and Schmidt, (2016) introduces more adaptive shrinkage via a local-global hierarchy. Each $\beta_{j}$ is modeled as $\beta_{j}=B_{\mathrm{raw},j}\,\lambda_{j}\,\tau$, where $B_{\mathrm{raw},j}\sim\mathcal{N}(0,1)$, $\lambda_{j}\sim\mathrm{C}^{+}(0,1)$ (local scale), and $\tau\sim\mathrm{C}^{+}(0,1)$ (global scale). Small coefficients are heavily shrunk by small local scales, while the heavy-tailed Cauchy priors allow some large signals to remain.
@@
Finally, the Bayesian lasso Park and Casella, (2008) imposes a Laplace (double-exponential) prior,
@@ 2.2 Frequentist Methods
We use a James–Stein-like shrinkage approach for VAR coefficients Giannone et al., (2015); Del Negro and Giannoni, (2015), implemented in R via VARshrink with method="ns". Instead of explicitly solving (3), the NS method estimates the necessary sample covariances of $\mathbf{y}_{t-i},\mathbf{y}_{t}$} and then applies a closed-form shrinkage rule to these covariance estimates, thereby producing a shrunk solution for $\bf{B}$. In this paper, we rely on the default choice for the shrinkage parameter, which VARshrink selects via a moment-based (empirical Bayes) formula akin to Stein’s unbiased risk estimate.
@@ 3.1 Data-Generating Processes
Following the textbook discussion in Lütkepohl, (2005, Sec. 2.1), we scale the simulated companion matrix so that its spectral radius is comfortably inside the unit circle. Specifically, we divide the initial draw $A_{1}$ by $1.1\,\rho_{\max}$, where $\rho_{\max}$ is the largest eigenvalue (in modulus). This simple rescaling keeps all roots of the characteristic polynomial strictly below one—thereby ensuring covariance–stationarity—while preserving empirically plausible coefficient magnitudes for macro‑VAR applications.
@@ 4 Data Analysis
We illustrate our methods on the Canada dataset from the R package vars Pfaff, (2008), which provides quarterly macroeconomic observations on four Canadian variables spanning $T=84$ quarters (1980Q1–2000Q4): employment (e, in log-index form), productivity (prod, in log-index form measuring labor productivity), real wages (rw, in log-index form), and the unemployment rate (U, in percent). Economic considerations suggest these variables are jointly dependent, making a Vector Auto Regression (VAR)-based approach suitable.

### v3

@@ 1 Introduction
Vector autoregressive (VAR) models remain a cornerstone of multivariate time‑series analysis (Sims, 1980; Chan, 2020; Koop and Korobilis, 2013). A dd‑dimensional VAR($p$) posits that an observed series $\{\mathbf{y}_{t}\}_{t=1}^{T}$ satisfies
@@
where $\boldsymbol{\varepsilon}_{t}\!\sim\!\mathcal{N}(\mathbf{0},\Sigma_{\varepsilon})$ is white noise. VARs are ubiquitous in macroeconomics and finance; see Stock and Watson (2002); Crump et al. (2021); Carriero et al. (2022); Zhou and Chan (2023) for recent surveys and methodological refinements.
@@
Because a VAR($p$) contains $d^{2}p$ coefficients, dimensionality rises quickly with either $d$ or $p$ (Bańbura et al., 2010; Koop, 2013; Korobilis and Pettenuzzo, 2019). Shrinkage—penalising or down‑weighting coefficients toward simpler structures—therefore plays a critical role in stabilising estimation and forecasts (Stock and Watson, 2002; Chan, 2021), especially when $d^{2}p\gg T$ (Bai et al., 2022).
@@
Bayesian shrinkage. Local–global priors such as the horseshoe (Carvalho et al., 2010) or Bayesian lasso (Park and Casella, 2008) shrink most coefficients strongly while allowing a few to escape toward their likelihood values. Variants include the spike‑and‑slab (George and McCulloch, 1997), hierarchical horseshoes (Makalic and Schmidt, 2016; Prüser, 2021), order‑invariant priors (Chan et al., 2024), dynamic sparsification for hybrid TVP‑VARs (Chan, 2023), and structured factor‑augmented horseshoes (Zhou and Chan, 2023). Semi‑global or block‑specific shrinkage extends the idea by letting related groups share a common scale parameter (Gruber and Kastner, 2025; Prüser and Blagov, 2022). Computational advances now make these priors feasible in very large systems: fast variational Bayes (Bernardi et al., 2024), approximate Bayes for huge multi‑country VARs (Huber et al., 2022), and conjugate subspace shrinkage (Huber and Koop, 2023). Pandemic‑robust “outlier‑aware” priors improve real‑time performance during extreme events (Cascaldi-Garcia, 2022). Other recent Bayesian shrinkage methods in time series have been explored (Huber and Koop, 2023; Kowal et al., 2019; Katz et al., 2024, 2025).
@@
Frequentist and machine‑learning regularisation. Ridge regression (Doan et al., 1984) remains popular, while equation‑by‑equation Lasso or elastic‑net penalties (Tibshirani, 1996; Zou and Hastie, 2005; Sánchez García and Cruz Rambaud, 2022) yield sparse high‑dimensional VARs with frequentist guarantees under weak dependence (Masini et al., 2022). Envelope and reduced‑rank ideas provide additional dimension‑reduction tools (Samadi and Herath, 2023; Cubadda and Hecq, 2022), and non‑linear factor compression has recently been explored (Klieber, 2024). Structured penalties—group Lasso, hierarchical lag shrinkage, SCAD, MCP—continue to improve point forecasts and interpretability (Nicholson et al., 2017; Basu et al., 2019; Nicholson, 2020; Song and Bickel, 2019; Chen and Chen, 2021). New inferential results ensure valid impulse‑response analysis even after sparse estimation (Krampe et al., 2023).
@@
Forecasting and empirical performance. Time‑varying parameter BVARs with horseshoe or normal‑gamma shrinkage adapt well to structural change (Bitto and Frühwirth-Schnatter, 2019; Feldkircher et al., 2024). Factor‑adjusted network VARs (FNETS) separate common and idiosyncratic dynamics to enhance both forecasts and network interpretation (Barigozzi et al., 2024). Large‑scale empirical comparisons consistently find that local–global or semi‑global priors outperform purely global shrinkage and many frequentist alternatives in density and point forecasts (Huber and Feldkircher, 2019; Aprigliano, 2020; Gefang et al., 2023).
@@ 2.1 Bayesian Shrinkage Priors
where $\mathbf{L}_{\Omega}$ is the Cholesky factor of a correlation matrix with an LKJ prior (Lewandowski et al., 2009), and each component of $\sigma$ follows a half-Cauchy prior. This flexible structure permits correlation among the $d$ error components.
@@
The horseshoe prior (Carvalho et al., 2010; Makalic and Schmidt, 2016) introduces more adaptive shrinkage via a local-global hierarchy. Each $\beta_{j}$ is modeled as $\beta_{j}=B_{\mathrm{raw},j}\,\lambda_{j}\,\tau$, where $B_{\mathrm{raw},j}\sim\mathcal{N}(0,1)$, $\lambda_{j}\sim\mathrm{C}^{+}(0,1)$ (local scale), and $\tau\sim\mathrm{C}^{+}(0,1)$ (global scale). Small coefficients are heavily shrunk by small local scales, while the heavy-tailed Cauchy priors allow some large signals to remain.
@@
Finally, the Bayesian lasso (Park and Casella, 2008) imposes a Laplace (double-exponential) prior,
@@ 2.2 Frequentist Methods
We use a James–Stein-like shrinkage approach for VAR coefficients (Giannone et al., 2015; Del Negro and Giannoni, 2015), implemented in R via VARshrink with method="ns". Instead of explicitly solving (3), the NS method estimates the necessary sample covariances of $\mathbf{y}_{t-i},\mathbf{y}_{t}$} and then applies a closed-form shrinkage rule to these covariance estimates, thereby producing a shrunk solution for $\bf{B}$. In this paper, we rely on the default choice for the shrinkage parameter, which VARshrink selects via a moment-based (empirical Bayes) formula akin to Stein’s unbiased risk estimate.
@@ 3.1 Data-Generating Processes
Following the textbook discussion in Lütkepohl (2005, Sec. 2.1), we scale the simulated companion matrix so that its spectral radius is comfortably inside the unit circle. Specifically, we divide the initial draw $A_{1}$ by $1.1\,\rho_{\max}$, where $\rho_{\max}$ is the largest eigenvalue (in modulus). This simple rescaling keeps all roots of the characteristic polynomial strictly below one—thereby ensuring covariance–stationarity—while preserving empirically plausible coefficient magnitudes for macro‑VAR applications.
@@ 4 Data Analysis
We illustrate our methods on the Canada dataset from the R package vars (Pfaff, 2008), which provides quarterly macroeconomic observations on four Canadian variables spanning $T=84$ quarters (1980Q1–2000Q4): employment (e, in log-index form), productivity (prod, in log-index form measuring labor productivity), real wages (rw, in log-index form), and the unemployment rate (U, in percent). Economic considerations suggest these variables are jointly dependent, making a Vector Auto Regression (VAR)-based approach suitable.